OpenStudy (anonymous):

What is the solution to the system of equations x+3y=15 and 4x+2y=30

OpenStudy (anonymous):

@VCabral1134 @lyssa577 @Ashleyrh1999

OpenStudy (texaschic101):

x + 3y = 15 x = -3y + 15 so sub in -3y + 15 in for x in the other equation. 4x + 2y = 30 4(-3y + 15) + 2y = 30 --- distribute through the parenthesis -12y + 60 + 2y = 30 -- simplify -10y + 60 = 30 --- subtract 60 from both sides -10y = 30 - 60 -10y = -30 -- now divide both sides by -10 y = 3 x + 3y = 15....now sub in 3 for y and find x x + 3(3) = 15 x + 9 = 15 -- subtract 9 from both sides x = 15 - 9 x = 6 now we can check our answers... 4x + 2y = 30 4(6) + 2(3) = 30 24 + 6 = 30 30 = 30 (correct) so x = 6 and y = 3